There are two questions I need help with. Hope you can help, and thank you!!
1.The purity of gold is measured in karats, with pure gold being 24 karats. Other purities of gold are expressed as proportional parst of pure gold. Thus, 18 karat gold is 18/24, or 75% pure gold, 12 karat gold is 12/24, or 50%; and so on. How much 12 karat gold should be mixed with pure gold to obtain 60 grams of 16 karat gold?
2. An air rescue plane averages 300 miles per hour in still air. It carries fuel for 5 hours of flying time. If, upon takeoff, it encouters a wind of 30 miles per hour, how far can it fly and return safely? (assume that the wind remains constant)
: There are two questions I need help with. Hope you can help, and thank you!!
1.The purity of gold is measured in karats, with pure gold being 24 karats. Other purities of gold are expressed as proportional parst of pure gold. Thus, 18 karat gold is 18/24, or 75% pure gold, 12 karat gold is 12/24, or 50%; and so on. How much 12 karat gold should be mixed with pure gold to obtain 60 grams of 16 karat gold?
2. An air rescue plane averages 300 miles per hour in still air. It carries fuel for 5 hours of flying time. If, upon takeoff, it encouters a wind of 30 miles per hour, how far can it fly and return safely? (assume that the wind remains constant)

Answer by stanbon(18792)

The 1st question was answered earlier this evening. Can you find it?
-------------------
2. An air rescue plane averages 300 miles per hour in still air. It carries fuel for 5 hours of flying time. If, upon takeoff, it encounters a wind of 30 miles per hour, how far can it fly and return safely? (assume that the wind remains constant)
---------------
Original rate: 300 mph ; time = 5 hrs. ; distance = 1500 miles
--------------------
With 30 mph wind
rate = 330 mph ; time = x hrs; distance = 330x miles
-------
against 30 mph wind
rate = 270 mph ; time = 5-x hrs; distance = 270(5-x)= 1350-270x miles
---------------
EQUATION:
distance out = distance back
330x = 1350-270x
600x = 1350
x = 2.25 hrs (time with the wind)
Distance with the wind = 330*2.25 = 742.5
-------------
Round trip distance: 2*742.5 = 1485 miles
===============================
Cheers,
Stan H.